Using prices in 2020 as a base year, calculate real GDP in 2020 and 2021.


Q 2020 P 2020 Q 2021 P 2021


Apples 120 1 240 3


oranges 65 2 90 4


plams 80 4 120 5


(Q) represent Quantity and (P) price


1
Expert's answer
2021-01-30T16:49:32-0500

Let's calculate the real GDP in 2020 year:


"Real\ GDP_{2020}=120\cdot\$1+65\cdot\$2+80\cdot\$4=\$570."

Let's calculate the real GDP in 2021 year:


"Real\ GDP_{2021}=240\cdot\$1+90\cdot\$2+120\cdot\$4=\$900."